OpenStudy (anonymous):

How would you find the surface area of a sphere if the circumference is 3pi

myininaya (myininaya):

circumference=C=2pi*r and surface area=SA=4pi*r^2. Since SA=2*C*r, all we need to know is r. And since C=3pi=2pi*r then r=3/2.

myininaya (myininaya):

so SA=2*C*r=2*3pi*3/2=9pi

OpenStudy (anonymous):

huh

myininaya (myininaya):

huh what?

myininaya (myininaya):

can you plug the circumference into the surface area formula?

myininaya (myininaya):

yes